OpenStudy (anonymous):

solve for x: 4root over 2x-1=1 can anyone help me plzz!

OpenStudy (anonymous):

\[4\sqrt{2x-1}=1\] \[\sqrt{2x-1}=1/4\] square on both sides \[2x-1=1/16\] \[2x=1/16+1\] \[2x=17/16\] \[x=17/(2)(16)\] \[x=17/32\]

OpenStudy (anonymous):

got it jinausa

OpenStudy (anonymous):

thnkuu so much!!!